OpenStudy (anonymous):

Greg combined different weights of fertilizer, soil, and compost to make 29 kilograms (kg) of potting soil for some plants. The weight of fertilizer was x kg. The weight of soil was 2 kg more than three times the weight of fertilizer, and the weight of compost was half the weight of fertilizer. This is represented in the equation below: x + (3x + 2) + one half x = 29 What was the difference in the weights of soil and compost Greg combined? 3 kg 6 kg 14 kg 17 kg

OpenStudy (anonymous):

@Compassionate @wio @Whitemonsterbunny17

OpenStudy (anonymous):

@johnweldon1993 can you please help me with this question when u get the chance

OpenStudy (anonymous):

i think it's 17kg but im not real sure if it correct

OpenStudy (johnweldon1993):

Well lets check it We just solve this for 'x' x + (3x + 2) + one half x = 29 4x + 2 + 1/2x = 29 (9/2)x = 27 x = 6 Alright...so we know the soil part was (3x + 2) when we plug in 6 for 'x' we get 20 The compost part is half the fertilizer...well since the fertilizer was just 'x'....we know that is 6 and so half of 6 is 3 So yes...17 is indeed correct^ :)

OpenStudy (anonymous):

omg thank you so so so much your are such a great help ^-^

OpenStudy (johnweldon1993):

Haha well you did all the work ^_^ :) anytime hun

OpenStudy (anonymous):

^-^ do u mind helping me with few more im not real good math lol

OpenStudy (anonymous):

Ricky walked away from a mall at a constant velocity for the first 30 seconds. He stopped at an ice cream parlor for 40 seconds. He then turned around and walked toward the mall again at a constant velocity for 10 seconds. He met a friend, and they both walked away from the mall at a constant velocity for the next 20 seconds. Which of the following graphs best represents Ricky's motion? First straight line joins ordered pairs 0, 0 and 30, 20, second straight line joins 30, 20 and 70, 20, third straight line joins 70, 20 and 80, 30, and fourth straight line joins 80, 30 and 100, 0 First straight line joins ordered pairs 0, 0 and 30, 20, second straight line joins 30, 20 and 40, 5, third straight line joins 40, 5 and 60, 50 First straight line joins ordered pairs 0, 0 and 30, 20, second straight line joins 30, 20 and 70, 20, third straight line joins 70, 20 and 80, 5, and fourth straight line joins 80, 5 and 100, 50 First straight line joins ordered pairs 0, 0 and 30, 20, second straight line joins 30, 20 and 70, 20, third straight line joins 70, 20 and 80, 5, and fourth straight line joins 80, 5 and 90, 50

OpenStudy (anonymous):

i think it's "C"

OpenStudy (johnweldon1993):

Ready for this \[\large \color \red {\text{YOU ARE CORRECT!!}}\] lol :)

OpenStudy (anonymous):

lol yayay!! :D:D thank you ^-^

OpenStudy (johnweldon1993):

haha thank you for being smart already :D :P

OpenStudy (anonymous):

^-^ thank you lol for smart as well lol :D

OpenStudy (johnweldon1993):

darn notifications not telling me you replied! lol....well hey, its my job ^_^ lol no just kidding lol, but being an engineering student in a good university means I have to be good in math , and physics, but thats not needed here :P